Searches related to Given that a monetarist predicts velocity to be 5, graph the aggregate demand curve that results if the money supply is Kes 400 billion. If the money supply falls to Kes 50 billion. Explain the effect of this fall on the position of the aggregate demand curve.

1
Expert's answer
2021-01-27T07:22:32-0500

Answer:-

"Velocity = GDP \/money supply"

"5=GDP \/400"

"GDP = 400*5 =2000"

If fall in money supply by 50

"GDP = 350 *5 = 1750"

According to the question



AD shift left when total consumer spending decline

1)Consumers might spend less because the cost of living is rising or government taxes have increased.

2) Spend less- saving more because expectation prices rise in the future

3) Exchange rate may be high

4) Income may be decrease

GDP =AD in long term (according to the question,

Due to monetary affect money supply become less so that "AD" curve shift leftward.